For an extremely relativistic particle of rest energy e=mc² ,show that the momentum (P) is given by;
P=E[1-½(E/E)²] 
1
Expert's answer
2021-03-03T07:55:50-0500

Relativistic momentum p is classical momentum multiplied by the relativistic factor "\u03b3."

"p=\u03b3mu,"

where m is the rest mass of the object, "u" is its velocity relative to an observer, and the relativistic factor.

"γ= \large\frac{1}{\sqrt{1-\large\frac{u^2}{c^2}}}" "p=\large\frac{mu}{\sqrt{1-\large\frac{u^2}{c^2}}}" "E = mc^2 \to m = \large\frac{E}{c^2}"

is the three dimensional relativistic momentum of the object in the lab frame with magnitude |p| = p. The relativistic energy E and momentum p include the Lorentz factor defined by:

"{\displaystyle \gamma _{(\mathbf {u} )}={\frac {1}{\sqrt {1-{\frac {\mathbf {u} \cdot \mathbf {u} }{c^{2}}}}}}={\frac {1}{\sqrt {1-\left({\frac {u}{c}}\right)^{2}}}}}"

Some authors use relativistic mass defined by:

"{\displaystyle m=\gamma _{(\mathbf {u} )}m_{0}}"

although rest mass m0 has a more fundamental significance and will be used primarily over relativistic mass m in this article.

Squaring the 3-momentum gives:

"{\displaystyle p^{2}=\mathbf {p} \cdot \mathbf {p} ={\frac {m_{0}^{2}\mathbf {u} \cdot \mathbf {u} }{1-{\frac {\mathbf {u} \cdot \mathbf {u} }{c^{2}}}}}={\frac {m_{0}^{2}u^{2}}{1-\left({\frac {u}{c}}\right)^{2}}}}"

then solving for u2 and substituting into the Lorentz factor obtains its alternative form in terms of 3-momentum and mass, rather than 3-velocity:

"\gamma = \sqrt{1 + \left(\frac{p}{m_0 c}\right)^2}"

Inserting this form of the Lorentz factor into the energy equation:

"E = m_0c^2\sqrt{1 + \left(\frac{p}{m_0 c}\right)^2}" "\large(\frac{E}{E_o})^2" "=1+\large\frac{p^2}{E_o^2} \to" "p^2 = E_o^2 (E^2-E_o^2) \to p = E_o\sqrt{E^2-E_o^2}"